What is 5 over 6 plus 7 over 9
Answer by stanbon(75887)

What is 5 over 6 plus 7 over 9
------
(5/6) + (7/9) = (5*9+6*7)/(6*9) = (45+42)/54 = 87/54 = 1.611..
-------------
Cheers,
Stan H.
----------------